I approached it the same way. 1/3 chance of getting 1,000 * 1/2 chance of opening one of the other two boxes.thebigkats wrote:Hi:
I used simple probability to arrive the answer. Let me know if this approach is wrong and I just got lucky:
We absolutely need $1000 box to be open. Chances of that key getting into the right box is 1/3 (it could go into any box)
Assuming that this is actually in the right box, chances of other keys (2 left) in the right box (either of the two) is also 1/2 (only two possibilities - either both keys in right boxes or both in wrong boxes).
so to arrive >$1000 means that our possibility is 1/3 * 1/2 = 1/6
Certainly I could have gone down the path of listing out the combinations etc. but I got to this answer in few sec through probability
GMATLiveTeachStarts Nov 22
LIVE ONLINE CLASSES
Get Ready for GMAT Test Day Faster with Live Online Classes
with Josh Braslow, 20+ years in test prep
- Nov 22 · Josh Braslow
- Sun · 9:30 AM to 12:30 PM ET
- Included
- 40 hours of live online classes + 6 months of TTP OnDemand
- Attend the first class for free
- Every class is recorded, so you never fall behind













